### v1

@@ 3.1 Localized LIS method
Note $x_{\bot}$ and $x_{r}$ are independent under $\pi_{0}$, so that $\pi_{0}(x_{\bot}|x_{r})=\pi_{0}(x_{\bot})$. Also note in the transformed representation, the approximation takes the same form,
@@ 3.2 Approximation error
  • $\pi_{0}(x),\pi(x)$ have dependence graph $G$ that is $s$-sparse with ambient dimension $r$.
@@
Assume the target distribution $\pi$ satisfies 3.1. Consider its LLIS approximation $\pi_{r}$ (3.3). Then there exists a dimension independent constant $C_{\pi}$ s.t. the marginal error of the LLIS approximation satisfies
$\max_{j}W_{1}(\pi_{j},\pi_{r,j})\leq C_{\pi}\max_{k\in V}\left(\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k})+\sum_{j\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j})\right)^{1/2},$ (3.6)
@@
  • The expectation in (3.7) is taken w.r.t. the approximate distribution $\tilde{\pi}_{r}$, while in (3.2) it is taken w.r.t. the target distribution $\tilde{\pi}$. So that 3.2 is a posteriori error estimate instead of a priori estimate.
@@
(2) We would also like to comment that the summation $\sum_{j\in V}$ in (3.6) is still a dimension independent term. Since the second order derivatives ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})$ decay fast with $d_{G}(j,k)$, which is a result of the sparsity graphical assumption.
@@ 4 Application II: localized score matching
The score function of a distribution $p(x)$ is defined as $s(x):=\nabla\log p(x)$. In score matching [14, 21], one uses variational method to approximate $p(x)$ by fitting the score function. Given a hypothesis class $\mathscr{P}=\{p_{\theta}(x):\theta\in\Theta\}$ of distributions parameterized by $\theta$, the score matching aims to minimize the Fisher divergence over $\theta$:
$D_{\rm Fisher}(p_{\theta},p):=\mathbb{E}_{X\sim p}\left|s_{\theta}(X)-s(X)\right|^{2},\quad s_{\theta}(x):=\nabla\log p_{\theta}(x).$
Though the Fisher divergence is not directly computable, a simple integration by parts gives a tractable loss function
$J(\theta)=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X)\right)+\left|s_{\theta}(X)\right|^{2}\right].$
Here $J(\theta)=D_{\rm Fisher}(p_{\theta},p)-\mathbb{E}_{X\sim p}\left|\nabla\log p(X)\right|^{2}$, so that minimizing $J(\theta)$ is equivalent to minimizing the Fisher divergence. In practice, $J(\theta)$ is approximated by its empirical version: given data $(X_{1},\dots,X_{N})$ drawn from $p$, the empirical loss is
$\hat{J}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X_{i})\right)+\left|s_{\theta}(X_{i})\right|^{2}\right].$
The hypothesis class $\mathscr{P}$ is constructed such that the gradient of the score function $s_{\theta}(x)$ is easy to compute and optimize over, so that an approximate distribution $\hat{p}_{\theta}$ can be obtained efficiently by minimizing $\hat{J}(\theta)$.
Despite its simple implementation, fitting a generic score function in high-dimensional space is still computationally expensive. We propose to reduce the complexity by leveraging the localized structure, i.e. the target distribution is or can be approximated by a sparse graphical model. The locality in the model largely reduces the complexity of the score function, based on which we propose the localized score matching method.
@@ 4.1 Localized score matching
Factorization of sparse graphical models
@@
The key motivation of localized score matching is the existence of local factorization of the log density for the sprase graphical models. Consider a distribution $p(x)$ with dependence graph $G=(V,E)$. By Hammersley-Clifford theorem [16], one can factorize
$\log p(x)=\sum_{j\in V}u_{j}(x_{N_{j}}).$
Note $u_{j}$ is a function of $x_{N_{j}}\in\mathbb{R}^{d_{N_{j}}}$, where we denote the local dimension
$d_{N_{j}}:=\sum_{k\in N_{j}}d_{k}.$
Suppose $G$ is $s$-sparse in the sense of 2.5, then the maximal local dimension satisfies
$d_{\rm loc}:=\max_{j}d_{N_{j}}\leq(1+s)\max_{j}d_{j}.$ (4.1)
@@
In a sparse graphical model, the local dimension $d_{\rm loc}\ll d$, so that the complexity of $\log p(x)$ would be greatly reduced. This motivates to fit the target using distributions from a hypothesis class $\mathscr{P}_{G}$ that captures the localized structure. Specifically, we consider the hypothesis class
$\mathscr{P}_{G}=\left\{p_{\theta}(x)=\exp\left(\sum_{j\in V}u_{\theta,j}(x_{N_{j}})\right):u_{\theta,j}\in\mathcal{U}_{\theta,j}\right\},$
where $\mathcal{U}_{\theta,j}$ is some class of functions in $\mathbb{R}^{d_{N_{j}}}$ to be determined. The distributions in $\mathscr{P}_{G}$ have dependence graph $G$ by definition. Notice the size of the hypothesis space, measured by its metric entropy [24], is $O(\text{poly}(d)\exp(Cd_{\rm loc}))$, thus avoiding the curse of dimensionality.
Here we assume that the $G$ is prefixed based on prior knowledge of the target distribution. One can also learn the dependence graph from data, but we do not consider it here.
Localized loss function
@@
We define the local score function as $s_{j}(x):=\nabla_{j}\log p(x)$. To fit the local score function, it suffices to minimize the local loss
$J_{j}(\theta):=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X)\right)+\left|s_{\theta,j}(X)\right|^{2}\right],$ (4.2)
where $s_{\theta,j}(x)=\nabla_{j}\log p_{\theta}(x)$. Note for $p_{\theta}\in\mathscr{P}_{G}$, the local score $s_{\theta,j}$ can be computed locally by
$s_{\theta,j}(x)=\sum_{k\in N_{j}}\nabla_{j}u_{\theta,k}(x_{N_{k}}),$
which only concerns a small group of the local scores. To fit the whole score function, one could consider to minimize $\max_{v\in V}J_{j}(\theta)$. However, it is no longer equivalent to minimizing
$\max_{j\in V}\mathbb{E}_{X\sim p}\left|s_{\theta,j}(X)-s_{j}(X)\right|^{2}=\max_{j\in V}\left(J_{j}(\theta)+C_{j}\right),\quad C_{j}:=\mathbb{E}_{X\sim p}|s_{j}(X)|^{2}.$
The reason is that $C_{j}$ variates with $j$ and directly minimizing $\max_{j}J_{j}(\theta)$ may ignores those $j$ with larger $C_{j}$. Nevertheless, one can relax the problem to be a saddle point problem:
@@
Here $\Theta$ denotes the parameter space of the hypothesis class. Though $C_{j}$ is not directly accessible, one can replace it by some accessible lower bound of it in implementation. One also needs to replace the loss by the empirical version
$\begin{split}&\hat{J}_{\rm loc}(\theta)=\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(\hat{J}_{j}(\theta)+\lambda_{j}\right),\\ &\hat{J}_{j}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X_{i})\right)+\left|s_{\theta,j}(X_{i})\right|^{2}\right],\end{split}$ (4.4)
where $X_{1},\dots,X_{N}$ are i.i.d. samples from $p$.
4.2 Statistical error
In this section, we aim to quantify rate the convergence of the minimizer $\hat{\theta}_{*}$ of the empirical loss (4.4) to the minimizer $\theta_{*}$ of the population loss (4.3), i.e.
$\theta_{*}=\mathop{\arg\min}_{\theta\in\Theta}J_{\rm loc}(\theta),\quad\hat{\theta}_{*}=\mathop{\arg\min}_{\theta\in\Theta}\hat{J}_{\rm loc}(\theta)$
in terms of the number of samples $N$ and the full dimension $d$. Take the hypothesis class $\mathscr{P}$ larger to include the true distribution, i.e. $p\in\mathscr{P}$, so that $p_{\theta_{*}}=p$. Then such convergence rate quantifies the sample complexity to learn the target distribution, i.e. how many samples are needed to approximate the target distribution within a certain error.
For simplicity, we consider the target distribution $p(x)$ supported in the unit interval $\mathbb{I}^{d}=[0,1]^{d}$, and take the following hypothesis class
$\begin{split}\mathscr{P}=\Big{\{}p_{\theta}(x)=\exp&\left(\sum_{j\in V}u_{\theta,j}(x_{N_{j}})\right):mI\preceq-\nabla^{2}\log p_{\theta}\preceq MI,~{}u_{\theta,j}\in\mathcal{U}_{\theta,j}\Big{\}},\\ &\mathcal{U}_{\theta,j}=\left\{u_{\theta,j}\in C^{2}_{0}(\mathbb{I}^{d_{N_{j}}}):\|u_{\theta,j}\|_{C^{2}}\leq R\right\}.\end{split}$ (4.5)
In practice, it can be implemented, for instance, by the Input Convex Neural Networks [1].
@@ 4.2 Statistical error
Given a $s$-sparse graph $G=(V,E)$ with ambient dimension $r$. Consider approximating the target distribution $p(x)\in\mathcal{P}_{1}(\mathbb{I}^{d})$ using localized score matching with the hypothesis class $\mathscr{P}$ defined above. Then with probability at least $1-\delta$, it holds that
$\max_{v\in V}W_{1}(p_{j},\hat{p}_{j})\leq CRN^{-1/4}\left(\log\delta^{-1}+\log b\right)^{1/4},$
where $\hat{p}:=p_{\hat{\theta}_{*}}$, $C$ is a constant depending only on $s,r,m,M$ and the local dimension $d_{\rm loc}$, thus independent of the total dimension $d$; and $b$ is the number of blocks, and is the only term of $O(d)$.
@@
Proofs can be found in Appendix C. The theorem implies that sample complexity of the localized score mathching is bounded by $N_{\varepsilon}=O(\varepsilon^{-4}\log(d/\delta))$, with only logarithm dependence on the total dimension $d$, exactly characterizing that learning a localized distribution does not suffer from the curse of dimensionality. We also comment that this result does not achieve a minimax rate for estimating the target distribution $p$, as this is a corollary of the statistical estimation of the local score functions.
5 Application III: localized neural network surrogate
In this section, we consider the problem of constructing a surrogate for high dimensional distributions. It is ubiquitous that the target distribution is accessible, but is computationally expensive to evaluate. A common strategy is to construct a surrogate of the target distribution that is more computationally efficient, but still preserves the required accuracy. In this section, we investigate how the localized structure in the distribution helps to reduce the complexity of finding the surrogate.
As in the previous section, we assume the target distribution $p(x)$ has dependence graph $G=(V,E)$ that is $s$-sparse with ambient dimension $r$. We will construct a surrogate distribution $p_{\theta}(x)$ that has the same dependence graph $G$, and is easier to compute and sample from. There are various ways to construct the surrogate, e.g. vartional autoencoder [15], or the localized score matching in the previous section, but we do not specify the training method here.
In this section, we focus on the approximation capacity of the surrogate, and specifically consider using neural networks as the surrogate. We take the hypothesis class to be
$\mathscr{P}_{\rm NN}=\Big{\{}p_{\theta}(x)=\exp\left(\sum_{j\in V}u_{\theta,j}(x_{N_{j}})\right):mI\preceq-\nabla^{2}\log p_{\theta}\preceq MI,~{}u_{\theta,j}\in\Phi_{d_{N_{j}}}(L,W)\Big{\}},$ (5.1)
where we denote the fully connected neural network function class [28]
$\Phi_{d_{0}}(L,W)=\{\mathcal{N}_{\theta}:\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}|\theta=\{\mathcal{W}^{(l)},b^{(l)}\}_{l=1}^{L}\in\Theta_{d_{0}}(L,W)\}.$
Here $L,W$ are the depth and width of the neural function
$\mathcal{N}_{\theta}(x)=(\mathcal{W}^{(L)}\sigma(\cdot)+b^{(L)})\circ\cdots\circ(\mathcal{W}^{(1)}x+b^{(1)}),$
where $\sigma$ is the activation function. $\Theta_{d_{0}}(L,W)$ denotes the admissible parameter class
$\Theta_{d_{0}}(L,W)=$$\Big{\{}\{\mathcal{W}^{(l)},b^{(l)}\}_{l=1}^{L}:\mathcal{W}^{(1)}\in\mathbb{R}^{W\times d_{0}},~{}\mathcal{W}^{(l)}\in\mathbb{R}^{W\times W}~{}(1<l<L),$
$~{}\mathcal{W}^{(L)}\in\mathbb{R}^{1\times W};b^{(l)}\in\mathbb{R}^{W}~{}(1\leq l<L),~{}b^{(L)}\in\mathbb{R}\Big{\}}.$
In this section, we take the activation function to be the ReLU function $\sigma(x)=\max\{0,x\}$.
Note the complexity of the neural network is greatly reduced due to the sparse graphical structure. We show that
Theorem 5.1.
Given a $s$-sparse graph $G=(V,E)$ with ambient dimension $r$. Consider approximating the target distribution $p\in\mathscr{P}$ (4.5) using the neural network surrogate $p_{\theta}\in\mathscr{P}_{\rm NN}$ (5.1). Then there exists constants $C_{1},C_{2}$ depending only on the local dimension $d_{\rm loc}$, such that if we take the hyperparameters $L=C_{1}L_{0}\log L_{0}$ and $W=C_{2}W_{0}\log W_{0}$ for the NN class $\mathscr{P}_{\rm NN}$, then $\exists p_{\theta}\in\mathscr{P}_{\rm NN}$ s.t.
$\max_{j\in V}W_{1}(p_{j},p_{\theta,j})\leq CRL_{0}^{-2/d_{\rm loc}}W_{0}^{-2/d_{\rm loc}}.$
Here $C$ is a constant depending on $s,r,m,M$ and the local dimension $d_{\rm loc}$, thus independent of the total dimension $d$.
Proofs can be found in Appendix D. Similar to 4.1, the above theorem shows in the approximation perspective that learning the localized distribution does not suffer from the curse of dimensionality.
6 Proof on the gradient estimate of Stein equation
In this section, we illustrate the main steps to obtain gradient estimate of the Stein equation, which is the core of 2.2 and 2.7.

### v2

@@ 4.2.2 Localized LIS Approximation
Note $\pi_{0}(x_{\bot}|x_{r})=\pi_{0}(x_{\bot})$ since $x_{\bot}$ and $x_{r}$ are independent under $\pi_{0}$. Also note in the whitened coordinate, the approximation takes the same form,
@@ 4.3 Approximation Error
  • $\pi_{0}(x),\pi(x)$ have dependency graph $G$ that is $(S,\nu)$-local.
@@
Assume the target distribution $\pi$ satisfies 4.1. Consider its LLIS approximation $\pi_{r}$ (4.3). Then there exists a dimension independent constant $C_{\pi}$ s.t. the marginal error of the LLIS approximation satisfies
$\max_{i}\mathsf{W}_{1}(\pi_{i},\pi_{r,i})\leq C_{\pi}\max_{k\in V}\left(\mathcal{R}(\widetilde{\mathcal{G}}_{k},\widetilde{\mathcal{X}}_{r,k})+\sum_{j\in V}\mathcal{R}(\widetilde{\mathcal{H}}_{j,k},\widetilde{\mathcal{X}}_{r,j})\right)^{1/2},$ (4.6)
@@
  • The expectation in (4.7) is taken w.r.t. the approximate distribution $\tilde{\pi}_{r}$, while in (4.2) it is taken w.r.t. the target distribution $\tilde{\pi}$. So that 4.1 is a posteriori error estimate instead of a priori estimate.
@@
(2) We would also like to comment that the summation $\sum_{j\in V}$ in (4.6) is still a dimension independent term. Since the second order derivatives ~$\tilde{\nabla}_{j}\tilde{\nabla}_{k}\log\tilde{l}(\tilde{x})$ decay exponentially with $\mathsf{d}_{G}(j,k)$, which is a result of the sparsity graphical assumption.
@@ 5 Application II: Localized Score Matching
5.1 Review of Score Matching
Score matching is a variational approach to approximating a probability distribution $p(x)$ by fitting its score function $s(x):=\nabla\log p(x)$. Originally introduced in [26] and later extended in [37], score matching circumvents the need for explicit density estimation by minimizing the Fisher divergence between a candidate distribution and the target distribution.
Given a parametric family of distributions $\mathscr{P}=\{p_{\theta}:\theta\in\Theta\}$, the score matching aims to find the parameter $\theta$ that minimizes the Fisher divergence
$\mathsf{F}(p\|p_{\theta}):=\mathbb{E}_{X\sim p}\left\|s_{\theta}(X)-s(X)\right\|^{2},\quad s_{\theta}(x):=\nabla\log p_{\theta}(x).$
While the Fisher divergence is generally not directly computable, an integration by parts yields a tractable objective function:
$J(\theta):=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X)\right)+\left\|s_{\theta}(X)\right\|^{2}\right].$
Note $J(\theta)$ only differs with the Fisher divergence $\mathsf{F}(p\|p_{\theta})$ by a constant:
$J(\theta)=\mathsf{F}(p\|p_{\theta})-\mathbb{E}_{X\sim p}\left\|\nabla\log p(X)\right\|^{2}.$
Thus minimizing $J(\theta)$ is equivalent to minimizing the Fisher divergence. In practice, given the data $(X_{1},\dots,X_{N})$ sampled from $p$, $J(\theta)$ is estimated by its empirical version:
$\hat{J}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla s_{\theta}(X_{i})\right)+\left\|s_{\theta}(X_{i})\right\|^{2}\right].$
The hypothesis class $\mathscr{P}$ is typically designed so that the score function $s_{\theta}(x)$ and its gradient are computationally tractable, enabling efficient optimization of $\hat{J}(\theta)$ and yielding an approximate distribution $\hat{p}_{\theta}$.
@@ 5.2 Localized Score Matching
Despite its simple implementation, score matching in high-dimensional spaces becomes computationally expensive due to the complexity of learning a generic score function. To mitigate this issue, we propose leveraging the locality structure in the target distribution. Specifically, if the distribution can be represented or well-approximated by a sparse graphical model, the score function should exhibit a lower complexity. Exploiting this locality, we introduce the localized score matching method, which significantly reduces computational costs while preserving approximation accuracy.
5.2.1 Clique Decomposition of Sparse Graphical Models
@@
The key motivation of localized score matching is the clique decomposition [4, 29] of the sprase graphical models, which a local factorization of the log density. Consider a distribution $p(x)$ with dependency graph $G=(V,E)$. By Hammersley-Clifford theorem [4], one can factorize
$\log p(x)=\sum_{j\in V}u_{j}(x_{\mathcal{N}_{j}}).$ (5.1)
Here $u_{j}$ is a function of $x_{\mathcal{N}_{j}}\in\mathbb{R}^{d_{\mathcal{N}_{j}}}$, and $d_{\mathcal{N}_{j}}$ denotes the local dimension
$d_{\mathcal{N}_{j}}:=\sum_{k\in\mathcal{N}_{j}}d_{k}.$
Suppose $G$ is $(S,\nu)$-local in the sense of 2.2, then the maximal local dimension satisfies
$d_{\rm loc}:=\max_{j}d_{\mathcal{N}_{j}}\leq(1+s)\max_{j}d_{j}.$ (5.2)
@@
Since $d_{\rm loc}\ll d$ for a sparse graphical model, the complexity of learning $\log p(x)$ is greatly reduced. This motivates us to approximate the target distribution using a hypothesis class $\mathscr{P}_{G}$ that explicitly incorporates the locality structure. Specifically, we define
$\mathscr{P}_{G}=\left\{p_{\theta}(x)=\exp\left(\sum_{j\in V}u_{\theta,j}(x_{\mathcal{N}_{j}})\right):u_{\theta,j}\in\mathcal{U}_{\theta,j}\right\},$
where $\mathcal{U}_{\theta,j}$ is a class of functions in $\mathbb{R}^{d_{\mathcal{N}_{j}}}$ to be specified. Distributions in $\mathscr{P}_{G}$ have dependency graph $G$ by construction. Notice the size of the hypothesis space, measured by its metric entropy [43], scales as $O(d^{\alpha}\exp(Cd_{\rm loc}))$, where $\alpha>0$ and $d_{\rm loc}$ is the maximal local dimensional. This structure effectively mitigates the curse of dimensionality, making learning and optimization more tractable in high-dimensional settings. Here we assume that the $G$ is prefixed based on prior knowledge of the target distribution. One can also learn the dependency graph from data, but we do not consider it here.
5.2.2 Localized Score Matching
By the clique decomposition (5.1), the score function of a MGF can be computed as
$s_{j}(x):=\nabla_{j}\log p(x)=\sum_{k\in\mathcal{N}_{j}}\nabla_{j}u_{k}(x_{\mathcal{N}_{k}}).$
@@ 5.2.2 Localized Score Matching
Here $s_{j}(x)$ is the $j$-th component of the score function, and we will refer to it as the local score function. To fit the local score, it suffices to minimize
$\mathbb{E}_{X\sim p}\left\|s_{\theta,j}(x)-s_{j}(x)\right\|^{2},$
where $s_{\theta,j}(x)=\nabla_{j}\log p_{\theta}(x)$ and $p_{\theta}\in\mathscr{P}_{G}$. By integrating by parts, it is equivalent to minimize
$J_{j}(\theta):=\mathbb{E}_{X\sim p}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X)\right)+\left\|s_{\theta,j}(X)\right\|^{2}\right],$ (5.3)
To estimate the full score function, one may consider minimizing $\max_{j\in V}J_{j}(\theta)$. However, this is not equivalent to minimizing
$\max_{j\in V}\mathbb{E}_{X\sim p}\left\|s_{\theta,j}(X)-s_{j}(X)\right\|^{2}=\max_{j\in V}\left(J_{j}(\theta)+C_{j}\right),\quad C_{j}:=\mathbb{E}_{X\sim p}\left\|s_{j}(X)\right\|^{2}.$
The discrepancy arises because $C_{j}$ varies with $j$ and directly minimizing $\max_{j}J_{j}(\theta)$ may neglect indices $j$ for which $C_{j}$ is large. Nevertheless, it can be relaxed to a saddle point problem:
@@
Here, $\Theta$ denotes the parameter space of the hypothesis class. Though $C_{j}$ is not directly accessible, it can be replaced in practice by a computable lower bound. In practice, given the data $(X_{1},\dots,X_{N})$ sampled from $p$, one computes the empirical loss
$\begin{split}&\hat{J}_{\rm loc}(\theta)=\max_{\lambda_{j}\leq C_{j}\atop\forall j\in V}\max_{j\in V}\left(\hat{J}_{j}(\theta)+\lambda_{j}\right),\\ &\hat{J}_{j}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\left[2\text{\rm tr}\left(\nabla_{j}s_{\theta,j}(X_{i})\right)+\left\|s_{\theta,j}(X_{i})\right\|^{2}\right].\end{split}$ (5.5)
Note $\hat{J}_{j}$ can be computed locally, in the sense that the local score $s_{\theta,j}$
$s_{\theta,j}(x)=\sum_{k\in\mathcal{N}_{j}}\nabla_{j}u_{\theta,k}(x_{\mathcal{N}_{k}})$
only concerns a small group of the local potential functions $u_{\theta,k}$. This facilitates efficient optimization of the localized score matching loss.
5.3 Reduction in Sample Complexity
In this section, we aim to show that the localized score matching reduces the sample complexity. To this end, we analyze the convergence rate of the empirical minimizer $\hat{\theta}_{*}$ of the empirical loss (5.5) to the population minimizer $\theta_{*}$ of the true loss (5.4), given by
$\theta_{*}=\mathop{\arg\min}_{\theta\in\Theta}J_{\rm loc}(\theta),\quad\hat{\theta}_{*}=\mathop{\arg\min}_{\theta\in\Theta}\hat{J}_{\rm loc}(\theta).$ (5.6)
Our goal is to quantify this convergence rate in terms of the number of samples $N$ and the full ambient dimension $d$. By enlarging the hypothesis class to include the target distribution, i.e. assuming $p\in\mathscr{P}$ so that $p_{\theta_{*}}=p$, this convergence rate provides an upper bound of the sample complexity of learning the target distribution, quantified by the number of samples required to approximate the target distribution within a given error.
For simplicity, we assume that the target distribution $p(x)$ is supported in the unit hypercube $\mathbb{I}^{d}=[0,1]^{d}$, and consider the hypothesis class
$\begin{split}\mathscr{P}=\Big{\{}p_{\theta}(x)=\exp&\left(\sum_{j\in V}u_{\theta,j}(x_{\mathcal{N}_{j}})\right):mI\preceq-\nabla^{2}\log p_{\theta}\preceq MI,~{}u_{\theta,j}\in\mathcal{U}_{\theta,j}\Big{\}},\\ &\mathcal{U}_{\theta,j}=\left\{u_{\theta,j}\in C^{2}_{0}(\mathbb{I}^{d_{\mathcal{N}_{j}}}):\|u_{\theta,j}\|_{C^{2}}\leq R\right\}.\end{split}$ (5.7)
In practice, this class can be implemented using Input Convex Neural Networks (ICNNs) [1].
@@ 5.3 Reduction in Sample Complexity
Let $G=(V,E)$ be a $(S,\nu)$-local graph. Consider approximating the target distribution $p(x)\in\mathcal{P}_{1}(\mathbb{I}^{d})$ using localized score matching with the hypothesis class defined in (5.7) and assume $p\in\mathscr{P}$. Then with probability at least $1-\lambda$, it holds that
$\max_{i}\mathsf{W}_{1}(p_{i},\hat{p}_{i})\leq CN^{-1/4}\left(\log\lambda^{-1}+\log b\right)^{1/4},$
where $\hat{p}:=p_{\hat{\theta}_{*}}$ (5.6) is the learned distribution. Here
  • $C$ is a constant depending only on $S,\nu,m,M,R$ and the local dimension $d_{\rm loc}$, but not on the total dimension $d$.
  • $b$ is the number of blocks, and is the only term that scales as $O(d)$.
@@
Proofs can be found in Appendix C. This result highlights a key advantage: despite the high ambient dimension $d$, the convergence rate depends only on the local dimension, making the method scalable in high dimensions. The theorem also implies that the sample complexity of localized score matching is bounded by $N_{\varepsilon}=O(\varepsilon^{-4}\log(d/\lambda))$, with only logarithm dependence on the total dimension $d$. This result precisely characterizes the fact that learning a localized distribution circumvents the curse of dimensionality.
6 Proofs
6.1 Explicit Solution of the Stein Equation
Lemma 6.1.
Under the conditions in 2.1 or 2.2, for any $1$-Lipschitz function $\phi:\mathbb{R}^{d_{i}}\to\mathbb{R}$, the solution of the marginal Stein equation (2.2) is given by (up to a constant)
$u(x)=-\int_{0}^{\infty}\mathbb{E}\left(\phi(X_{t,i}^{x})-\mathbb{E}_{\pi}[\phi(x_{i})]\right)\mathrm{d}t.$ (6.1)
where $X_{t}^{x}$ is the path solution of the overdamped Langevin dynamics
$\mathrm{d}X_{t}^{x}=\nabla\log\pi(X_{t}^{x})\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},\quad X_{0}^{x}=x.$ (6.2)
As a corollary, the gradient estimate holds
$\left\|\nabla_{j}u(x)\right\|\leq\int_{0}^{\infty}\mathbb{E}\|\nabla_{x_{j}}X_{t,i}^{x}\|\mathrm{d}t.$ (6.3)
Remark 6.1.
Note the generator of the Langevin dynamics (6.2) is $\mathcal{L}_{\pi}=\Delta+\nabla\log\pi\cdot\nabla$. Then the marginal Stein equation (2.2) can be written as
$\mathcal{L}_{\pi}u=\Delta u+\nabla\log\pi\cdot\nabla u=\phi(x_{i})-\mathbb{E}_{\pi}[\phi(x_{i})].$
This type of Stein equation is known to admit explicit solutions of the form (6.1) (see [2]).
Proof.
Let $X_{t}^{x}$ solves the Langevin dynamics (6.2). By Dynkin’s formula [33],
$\begin{split}\mathbb{E}u(X_{T}^{x})-u(x)=~{}&\mathbb{E}\int_{0}^{T}\left(\nabla\log\pi(X_{t}^{x})\cdot\nabla u(X_{t}^{x})+\Delta u(X_{t}^{x})\right)\mathrm{d}t\\ =~{}&\int_{0}^{T}\mathbb{E}\left(\phi(X_{t,i}^{x})-\mathbb{E}_{\pi}[\phi(x_{i})]\right)\mathrm{d}t.\end{split}$ (6.4)
Under the conditions in 2.1 or 2.2, the distribution is strongly log-concave (see 2.2), and it is well-known that Law$(X_{t}^{x})$ converges to the equilibrium $\pi$ exponentially. It implies that the limit $T\rightarrow\infty$ exists for both sides in (6.4), and the limit is
$\int u(x)\pi(x)\mathrm{d}x-u(x)=\int_{0}^{\infty}\mathbb{E}\left(\phi(X_{t,i}^{x})-\mathbb{E}_{\pi}[\phi(x_{i})]\right)\mathrm{d}t.$
This gives (6.1) up to a constant. Taking derivative w.r.t $x_{j}$ gives
$\nabla_{j}u(x)=-\int_{0}^{\infty}\mathbb{E}\left[\nabla_{x_{j}}X_{t,i}^{x}\cdot\nabla\phi(X_{t,i}^{x})\right]\mathrm{d}t.$
Note taking derivative on both sides is valid due to the exponential decay of $\nabla_{x_{j}}X_{t,i}^{x}$. Since $\phi$ is $1$-Lipschitz, we obtain
$\left\|\nabla_{x_{j}}u(x)\right\|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{x_{j}}X_{t,i}^{x}\|\left\|\nabla\phi(X_{t,i}^{x})\right\|\right]\mathrm{d}t\leq\int_{0}^{\infty}\mathbb{E}\|\nabla_{x_{j}}X_{t,i}^{x}\|\mathrm{d}t.$
This completes the proof. ∎
6.2 Proof of 2.1
Before the proof, we introduce the following technical lemma, which essentially controls the diffusion speed in the sparse graphs.
Lemma 6.2.
Let $H_{t}\in\mathbb{R}^{d\times d}$ be a time-dependent positive definite matrix satisfying:
  1. a.
    $H_{t}$ has dependency graph $G$, i.e. $H_{t}(i,j)=0$ if $\mathsf{d}_{G}(i,j)>1$.
  2. b.
    $\exists M>0$ s.t. $\forall t\geq 0,0\preceq H_{t}\preceq MI_{d}$.
Consider the matrix ODE
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=-H_{t}G_{t},\quad G_{0}=I_{d}.$ (6.5)
Then for any $t\geq 0$, it holds that
$\|G_{t}(i,j)\|\leq\exp(-tM)\sum_{k=\mathsf{d}_{G}(i,j)}^{\infty}\frac{t^{k}M^{k}}{k!}.$
Proofs can be found in Section A.1, which is based on the polynomial approximation trick in [14] and Dyson series [15].
Proof of 2.1.
Under the conditions in 2.1, 6.1 holds. By (6.3), it remains to control $\nabla_{x_{j}}X_{t,i}^{x}$. Taking derivative w.r.t. $x$ in (6.2), we get
Note here $\nabla X_{t}^{x}=\nabla_{x}X_{t}^{x}\in\mathbb{R}^{d\times d}$. Denote $G_{t}=\mathrm{e}^{mt}\nabla X_{t}^{x}$ and $\tilde{H}_{t}=H_{t}-mI$, then it holds that
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=\mathrm{e}^{mt}\left(m\nabla X_{t}^{x}-H_{t}\nabla X_{t}^{x}\right)=-\tilde{H}_{t}G_{t},\quad G_{0}=\nabla X_{0}^{x}=I_{d}.$
By assumption, $0\preceq\tilde{H}_{t}\preceq(M-m)I$, and $\tilde{H}_{t}(i,j)=0$ if $\mathsf{d}_{G}(i,j)>1$. By 6.2,
$\mathrm{e}^{mt}\|\nabla_{x_{j}}X_{t,k}^{x}\|=\|G_{t}(k,j)\|\leq\mathrm{e}^{-(M-m)t}\sum_{q=\mathsf{d}_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}.$
$\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\leq~{}$ $\mathrm{e}^{-mt}\mathrm{e}^{-(M-m)t}\sum_{j\in V}\sum_{q=\mathsf{d}_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}$
[…]
Here we use the sparsity condition $|\mathcal{N}_{j}^{q}\backslash\{v\}|\leq Sq^{\nu}$ (2.5). From (6.3), we obtain
$\sum_{j\in V}\|\nabla_{j}u(x^{(j)})\|\leq~{}$ $\sum_{j\in V}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t=\mathbb{E}\int_{0}^{\infty}\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t$
[…]
By A.2, it holds that (denote $\kappa=\frac{M}{m}$)
$\sum_{j\in V}\|\nabla_{j}u(x^{(j)})\|\leq~{}$ $\frac{1}{m}+\frac{S}{M}\nu!\left(\frac{m}{M}\right)^{-\nu-1}\left(1-\frac{m}{M}\right)$
[…]
Taking supremum over $x^{(j)}$, we obtain the desired gradient estimate with $\delta=\frac{S\nu!\kappa^{\nu}}{m}$. ∎